OpenStudy (magbak):

Please help!! Medal will be awarded. Can some one create a line of best fit for this graph.

OpenStudy (magbak):

y =1/2x y =-1/2x y = 2x y = –2x These are the options.

OpenStudy (anonymous):

The graph is going up, so the slope must be positive

OpenStudy (magbak):

Tell me is is A or is it C.

OpenStudy (anonymous):

Its c lol

OpenStudy (magbak):

Thanx

OpenStudy (anonymous):

np

OpenStudy (magbak):

Thank you.

OpenStudy (magbak):

Thank you.